# Collaborative Object Handover in a Robot Crafting Assistant

### v1

@@ V-A Handover episodes
As shown in Table II, over a third of handover episodes were rated as having poor quality by the observer (i.e., bad handovers). While the predictions are binary or 3-way classifications, the diversity and subtlety in how people signalled their intent during crafting made the task challenging. We reviewed the observation notes and identified two types of errors that caused a bad episode. One type is motion primitive error due to focusing the adaptation at key state transitions (see Figure 1). Specifically, once object transfer completion is predicted and the robot begins to execute action primitive 4, the autonomous model stops monitoring the participant’s behaviours. Thus, when the robot is tucking its arm, it is unable to wait or react if a participant signals to the robot at this time. The other type is prediction error due to imperfect recognition of participants’ true intentions. This can be further divided into incorrect prediction in handover timing, such as interrupting a participant’s crafting activities by starting an episode early, and incorrect prediction in handover location, such as colliding with a birdhouse kept in the middle of the work desk by performing a middle handover. Motion primitive error was the most common type in Stage 1 with the observer reporting 11 out of 20 participants who had at least one noticeable occurrence, while timing prediction error was the most common in Stages 2 (12 out of 20) and 3 (15 out of 20). The observation notes revealed that motion primitive error was mainly caused by participants attempting to perform bidirectional handovers instead of R2H or H2R handovers. Timing prediction error was mainly caused by the robot misrecognising a participant’s assembly or painting gestures as episode or object transfer initiations, thus moving towards the participant or stretching out its arm too early, interrupting the participant.
V-B Participant behaviour and perception
Participants showed diverse behaviours when communicating their intention to the robot with a combination of gaze and hand gestures, similar to observations in the teleoperated sessions [11]. Recall that participants were not told if the robot was fully autonomous or being teleoperated at the start of the experiment and were asked to guess this post-study. Three participants guessed the robot was teleoperated, four were unsure, and the remaining 13 guessed the robot was autonomous. In the interview, participants elaborated that the reasons for guessing the robot was teleoperated included the robot being responsive to gaze and demonstrating good collaboration that was considered to be on par with a human; The reason for being unsure was due to a mix of episodes of good collaboration and robot errors; The reasons for guessing the robot was autonomous were handover errors which they expected a human would be able to avoid.
@@ V-B Participant behaviour and perception
As for negative impressions, 10 participants reported handover timing errors, including six participants reporting feeling rushed or the robot being impatient at times, and two commenting that the robot interrupted or surprised them; three reported handover location errors, including the left-handed participant; three commented the robot was inconsistent or difficult to predict; three considered having the robot as an assistant to be less efficient than crafting by themselves or the robot was being too slow; two commented that while they noticed the robot’s adaptation it was not to a satisfactory performance. Compared to handover errors reported by the observer (Section V-A), fewer participants discussed robot errors in the interview. This is consistent with [11], where the participants’ focus on the overall crafting experience, as opposed to the operator’s focus on the individual handover episodes, led to differences in their perception of handover quality.
VI Teleoperated vs. Autonomous Handover
VI-A Handover episodes
We compared the handover episodes when the robot was teleoperated [11] and when the robot was autonomous. Note that the teleoperated csv data included all 20 participants, while handover csv data from P6, P10, P14, P16 in the autonomous sessions were removed (see Section V-A). We visualise the distribution of percentages of OTP location, handover type and quality, as well as handover timings when the robot was autonomous vs. teleoperated in Figure 2.
@@ VI-A Handover episodes
In terms of handover quality, the observer rated more teleoperated handovers as good (P(TeleAuto)=99.29%, 90% CI [0.06, 0.36]) or neutral (P(TeleAuto)=85.49%, 90% CI [-0.07, 0.38]), while more autonomous handovers were rated as bad (P(AutoTele)=100%, 90% CI [2.64, 4.83]).
A one-way ANOVA showed that the handover model (teleoperated or autonomous) had a significant influence on average episode duration (, , medium effect with ). The teleoperated episodes were longer on average, with shorter pauses at the workspace (, , large effect with ) and longer pauses at the storage space (, , large effect with ).
In this box plot, teleoperated sessions have bigger spread and higher mean in choosing left OTP, autonomous sessions have smaller spread and higher mean in choosing middle OTP, both sessions have very few right OTPs.
In this box plot, teleoperated sessions have lower mean in choosing R2H, autonomous sessions have higher mean in choosing H2R, teleoperated sessions have higher mean and wider spread in choosing bidirectional handovers.
In this box plot, teleoperated sessions have higher mean in good handovers, autonomous sessions have higher mean in bad handovers, autonomous sessions have lower mean and wider spread in neutral handovers.
In this box plot, teleoperated sessions have higher mean in episode duration and pause at storage space, autonomous sessions have higher mean in pause at work space.
VI-B Subjective perception
We report participants’ questionnaire responses in the teleoperated and autonomous sessions in Table III. Participants reported similar impressions of the robot, except for perceived intelligence where a Wilcoxon signed rank test showed significantly higher ratings for the teleoperated session (, ). Regarding HRC impressions, the teleoperated sessions have significantly higher ratings in fluency (, ), trust (, ), and working alliance (, ).
@@ VI-B Subjective perception
Further, when analysing subjective ratings for each of the three crafting stages, an increasing trend from Stage 1 to Stage 3 was found when the robot employed the autonomous handover policy, albeit not significant. When the robot was teleoperated, a significant increase in subjective perception from Stage 1 to Stage 3 was found [11]. We hypothesise that the significant increase in subjective perception when the robot was teleoperated was caused by both increased familiarity with the robot and the crafting task, as well as the human operator’s adaptation of the robot’s handovers to an individual participant’s working styles and preferences over time. As the autonomous robot used the same handover policy throughout the session without personalising to each individual participant, while participants had increased familiarity with the robot and the task, the lack of personalisation limited how much their perception would increase over time. We further analyse personalisation by the human operator in Section VI-C.
VI-C Personalisation in teleoperated sessions
We analysed the timing and location of teleoperated handovers across all 20 participants aligned by the episode ID. As shown in Fig 3a, in the Preparation stage the operator chose middle OTP for most participants, while a mix of OTPs were used later. This suggests that the operator chose middle OTP as a general strategy at the start of each session, and then personalised the OTPs as the session progressed. Further, we examined the accuracy of the autonomous model for predicting OTPs on the FACT HRC dataset using CV by episodes (see Section IV-A). As shown in Table IV, the model yielded the highest accuracy when predicting the first 20% of episodes in each participant (trained on the remaining 80% of episodes) with a declining accuracy in predicting OTPs for handovers that occurred later. The increased prediction errors in later episodes indicates that the autonomous policy diverged from the human operator’s behaviours more in later episodes. This supports our assumption that the operator began with a generic handover strategy but personalised it as the collaboration progressed.
@@ VI-C Personalisation in teleoperated sessions
The x-axis is episode number, the y-axis is OTP location choice (%). Episode 0 to 5 is approximately preparation stage, 5 to 15 is assemble stage, 15 to 25 is painting stage, with overlaps between the stages. The percentage of choice of middle handover is near 100% at the start, but drop to around 50% close to episode 10. The percentage of choice of left handover raised from close to 0% to around 50% close to episode 10. The percentage of right handovers remain low with small increases after episode 10.
The x-axis is episode number, the y-axis is std of pause with a value range of 0 to 400 seconds. Episode 0 to 5 is approximately preparation stage, 5 to 15 is assemble stage, 15 to 25 is painting stage, with overlaps between the stages. The std of pause at work space showed a peak at around episode 13 with value slightly above 100, and then at around episode 24 with value slightly above 200. The std of pause at storage space showed a peak at around episode 3 and then at around episode 13 with a value slightly above 100, with a tall peak at around episode 22 with value slightly above 350.
VII Discussion
@@
The machine learning experiments in Section IV illustrate that our autonomous handover model learned key temporal-spatial adaptation strategies adopted by a human operator to achieve collaborative human-robot handovers. However, the user study evaluation in Sections V & VI showed significant gaps between autonomous handovers and teleoperation, rejecting our hypothesis that the autonomous policy will achieve human-level collaboration performance.
@@
As discussed in [1], a handover episode can be divided into the pre-handover phase involving perception, planning and communication, and the physical handover phase involving controls and error handling. During the pre-handover phase, from the robot’s perspective, our analysis of handover errors in Sections V-A & VI-A showed that it is important to incorporate both human behaviours (hand gesture, head and body movements, gaze directions) and task context (task progress, object types and locations). From the user’s perspective, our analysis of user perceptions in Sections V-B & VI-B indicated that there may be a trade-off between reactive handovers (i.e., robot responding to explicit human cues) and proactive handovers (i.e., robot anticipating handovers based on past experiences and context understanding). Closer examination of the teleoperated handovers in Section VI-C showed that the human operator personalised their handover strategies for individual users. Therefore, it may be beneficial for autonomous handovers to incorporate prior knowledge during the pre-handover phase alongside observations specific to that episode, in order to allow correct perception of human states and intentions, as well as correct mapping from the perceived human state to their desirable handover policy (i.e., personalisation).
@@
During the physical handover phase, our analysis of autonomous handovers and participant behaviours in Sections V-A & V-B showed that autonomous handovers can benefit from increased transparency and explainability, allowing a user to understand the robot’s observations, plans, and actions and thus providing explicit or implicit feedback to adjust or correct them. This allows a human-in-the-loop approach to error handling and personalisation from the robot’s perspective, while increasing control and agency from the user’s perspective. However, such close human supervision may also conflict with the desire towards proactive handovers or efficient HRC. Thus, it is important to understand how different user and task context may influence the locus of control (robot vs. user) when designing collaborative handovers.
VIII Conclusions
@@
We developed an autonomous handover model using the FACT HRC dataset of teleoperated handovers [11], which predicts when and where a mobile manipulator robot performs key handover actions while engaging in task-oriented handovers. By conducting user study evaluation and comparing with teleoperated handovers, we identified open challenges in developing autonomous handovers that can achieve human-level HRC performance.

### v2

@@ 5.1 Handover episodes
As shown in Table 2, over a third of handover episodes were rated as having poor quality by the observer (i.e., bad handovers). While the predictions are binary or 3-way classifications, the diversity and subtlety in how people signalled their intent during crafting made the task challenging. We reviewed the observation notes and identified two types of errors that caused a bad episode. One type is motion primitive error due to focusing the adaptation at key state transitions (see Figure 1). Specifically, once object transfer completion is predicted and the robot begins to execute AP4, the autonomous model stops monitoring the participant’s behaviours. Thus, when the robot is tucking its arm, it is unable to wait or react if a participant signals to the robot at this time. The other type is prediction error due to imperfect recognition of participants’ true intentions. This can be further divided into incorrect prediction in handover timing, such as interrupting a participant’s crafting activities by starting an episode early, and incorrect prediction in handover location, such as colliding with a birdhouse kept in the middle of the work desk by performing a middle handover. Motion primitive error was the most common type in Stage 1 with the observer reporting 11 out of 20 participants who had at least one noticeable occurrence, while timing prediction error was the most common in Stages 2 (12 out of 20) and 3 (15 out of 20). The observation notes revealed that motion primitive error was mainly caused by participants attempting to perform bidirectional handovers instead of R2H or H2R handovers. Timing prediction error was mainly caused by the robot misrecognising a participant’s assembly or painting gestures as episode or object transfer initiations, thus moving towards the participant or stretching out its arm too early and interrupting them.
5.2 Participant behaviour and perception
Participants showed diverse behaviours when communicating their intention to the robot with a combination of gaze and hand gestures, similar to observations in the teleoperated sessions [?]. Recall that participants were not told if the robot was fully autonomous or being teleoperated at the start of the experiment and were asked to guess this post-study. Three participants guessed the robot was teleoperated, four were unsure, and the remaining 13 guessed correctly the robot was autonomous. In the interview, participants elaborated that the reasons for guessing the robot was teleoperated included the robot being responsive to gaze and demonstrating good collaboration that was considered to be on par with a human; The reason for being unsure was due to a mix of episodes of good collaboration and robot errors; The reasons for guessing the robot was autonomous were handover errors which they expected a human would be able to avoid.
@@ 5.2 Participant behaviour and perception
In terms of negative impressions, 10 participants reported handover timing errors, including six participants reporting feeling rushed or the robot being impatient at times, and two commenting that the robot interrupted or surprised them; three reported handover location errors, including the left-handed participant; three commented the robot was inconsistent or difficult to predict; three considered having the robot as an assistant to be less efficient than crafting by themselves or the robot was being too slow; two commented that while they noticed the robot’s adaptation it was not to a satisfactory performance. Compared to handover errors reported by the observer (see Section 5.1), fewer participants discussed robot errors in the interview. This is consistent with [?], where the participants’ focus on the overall crafting experience as a user, as opposed to the operator’s focus on the individual handover episodes as an observer, led to differences in their perception of handover quality.
6 Teleoperated vs. Auto Handover
6.1 Handover episodes
We compared the handover episodes when the robot was teleoperated [?] and when the robot was autonomous. Note that the teleoperated CSV data included all 20 participants, while handover CSV data from P6, P10, P14, P16 in the autonomous sessions were removed (see Section 5.1). We visualise the distribution of percentages of OTP location, handover type and quality, as well as handover timings when the robot was autonomous vs. teleoperated in Figure 2.
@@ 6.1 Handover episodes
Regarding handover quality, the observer rated more teleoperated handovers as good (P(TeleAuto)=99.29%, 90% CI [0.06, 0.36]) or neutral (P(TeleAuto)=85.49%, 90% CI [-0.07, 0.38]), while more auto handovers were rated as bad (P(AutoTele)=100%, 90% CI [2.64, 4.83]).
A one-way ANOVA showed that the handover model (teleoperated or autonomous) had a significant influence on average episode duration (, , medium effect with ). The teleoperated episodes were significantly longer on average, with shorter pauses at the workspace (, , large effect with ) and longer pauses at the storage space (, , large effect with ).
[…]
6.2 Subjective perception
We report participants’ questionnaire responses in the teleoperated and autonomous sessions in Table 3. Participants reported similar impressions of the robot, except for perceived intelligence where a Wilcoxon signed rank test showed significantly higher ratings for the teleoperated session (, ). Regarding HRC impressions, the teleoperated sessions have significantly higher ratings in fluency (, ), trust (, ), and working alliance (, ).
@@ 6.2 Subjective perception
Further, when analysing subjective ratings for each of the three crafting stages, an increasing trend from Stage 1 to Stage 3 was found when the robot employed the autonomous handover policy, albeit not significant. When the robot was teleoperated, a significant increase in subjective perception from Stage 1 to Stage 3 was found [?]. We hypothesise that the significant increase in subjective perception when the robot was teleoperated was caused by both increased familiarity with the robot and the crafting task, as well as the human operator’s adaptation of the robot’s handovers to an individual participant’s working styles and preferences over time. As the autonomous robot used the same handover policy throughout the session without personalising to each individual participant, while participants had increased familiarity with the robot and the task, the lack of personalisation limited how much their perception would increase over time. We further analyse personalisation by the human operator in Section 6.3.
6.3 Personalisation in teleoperated sessions
We analysed the timing and location of teleoperated handovers across all 20 participants aligned by the episode ID. As shown in Fig 3a, in the Preparation stage the operator chose middle OTP for most participants, while a mix of OTPs were used later. This suggests that the operator chose middle OTP as a general strategy at the start of a session then personalised the OTPs as the session progressed. Further, we examined the accuracy of the autonomous model for predicting OTPs on the FACT HRC dataset using CV by episodes (see Section 4.1). As shown in Table 4, the model yielded the highest accuracy when predicting the first 20% of episodes in each participant (trained on the remaining 80% of episodes) with a declining accuracy in predicting OTPs that occurred later. The increased prediction errors in later episodes indicates that the autonomous policy diverged from the human operator’s behaviours more in later episodes. This supports our assumption that the operator began with a generic handover strategy but personalised it as the collaboration progressed.
@@ 6.3 Personalisation in teleoperated sessions
[…]
7 Discussion
@@
The machine learning experiments in Section 4 illustrate that our autonomous handover model learned key temporal-spatial adaptation strategies adopted by a human operator to achieve collaborative human-robot handovers. However, the user study evaluation in Sections 5 & 6 showed significant gaps between autonomous handovers and teleoperation, rejecting our hypothesis that the autonomous policy will achieve human-level collaboration performance.
@@
As discussed in [?], a handover episode can be divided into the pre-handover phase involving perception, planning and communication, and the physical handover phase involving controls and error handling. During the pre-handover phase, from the robot’s perspective, our analysis of handover errors in Sections 5.1 & 6.1 showed that it is important to incorporate both human behaviours (hand gesture, head and body movements, gaze directions) and task context (task progress, object types and locations). From the user’s perspective, our analysis of user perceptions in Sections 5.2 & 6.2 indicated that there may be a trade-off between reactive handovers (i.e., robot responding to explicit human cues) and proactive handovers (i.e., robot anticipating handovers based on past experiences and context understanding). Closer examination of the teleoperated handovers in Section 6.3 showed that the human operator personalised their handover strategies for individual users. Therefore, it may be beneficial for autonomous handovers to incorporate prior knowledge during the pre-handover phase alongside observations specific to that episode, in order to allow correct perception of human states and intentions, as well as correct mapping from the perceived human state to their desirable handover policy (i.e., personalisation).
@@
During the physical handover phase, our analysis of autonomous handovers and participant behaviours in Sections 5.1 & 5.2 showed that autonomous handovers can benefit from increased transparency and explainability, allowing a user to understand the robot’s observations, plans, and actions and thus providing explicit or implicit feedback to adjust or correct them. This allows a human-in-the-loop approach to error handling and personalisation from the robot’s perspective, while increasing control and agency from the user’s perspective. However, such close human supervision may also conflict with the desire towards proactive handovers or efficient HRC with less user intervention. Thus, it is important to understand how different user and task context may influence the locus of control (robot vs. user) when designing collaborative handovers.
While providing insights on collaborative handovers, this study has a number of limitations. The simplification of the handover episode into a set of representative motion primitives (see Section 4.2) has limited the flexibility of the autonomous handover policy, especially in its ability to recover from recognition errors. An alternative to the supervised learning based approach for developing the autonomous handover policy may be behaviour cloning [?], which have shown benefits in learning complex and natural robot behaviours from human demonstrations. In addition, incorporating multimodal interaction such as speech [?] will increase the human-robot communication during handovers. Further, to generalise learnings from this study set in the crafting scenario to other HRC tasks, additional context should be considered, such as the human’s goals or potential risk of the activities. We recommend researchers to develop handover timing and location adaptation policies specific to their human, robot, and task context, as well as to incorporate suitable personalisation and locus of control to achieve collaborative handovers.
8 Conclusions
@@
We developed an autonomous handover model using the FACT HRC dataset of teleoperated handovers [?], which predicts when and where a mobile manipulator robot performs key handover actions while engaging in task-oriented handovers. By conducting user study evaluation and comparing with teleoperated handovers, we identified open challenges in developing autonomous handovers that can achieve human-level HRC performance.
Acknowledgments
This study was funded by the ARC Future Fellowship FT200100761.